OpenStudy (anonymous):

What is the energy of a photon of green light with a frequency of 5.96 ✕ 1014 s-1?

OpenStudy (anonymous):

use this equasion E=hf.

OpenStudy (anonymous):

equation*

OpenStudy (anonymous):

Can you help as to what I plug in for those variables?

OpenStudy (anonymous):

ok so h = Planck's constant = 6.63 x 10^-34 J s and your frequency is given and is equal to 5.96 ✕ 1014.now just miltiply.

OpenStudy (anonymous):

Would it be the same equation for "Calculate the energy of a photon of red light with a wavelength of 6.39 ✕ 10-5 cm." ?

OpenStudy (doc.brown):

Yes. Same formula. You may also use v to represent frequency, but it's the same math.\[E=hv\]

OpenStudy (doc.brown):

\[E=6.63\times 10^-34J\cdot s\times\frac{6.39\times 10^{-5}}{s}=4.24\times 10^{-38}J\]

OpenStudy (anonymous):

Oh! That makes sense! Thank you!!

OpenStudy (anonymous):

In the problem below, what does the s-1 mean? "Calculate the energy of a photon of green light with a frequency of 5.80 ✕ 1014 s-1"

OpenStudy (joannablackwelder):

s-1 means per second, as in waves per second.

OpenStudy (doc.brown):

\[x^{-n}=\frac{1}{x^n}\]so\[s^{-1}=\frac{1}{s}\]I'll just say, I use it exclusively. I never write m/s unless I'm explaining something to someone else. In my own work I write\[m\cdot s^{-1}\]